A spherical surface completely surrounds a collection of charges. Find the electric flux (with its sign) through the surface if the collection consists of (a) a single +2.50 × 10-6 C charge, (b) a single -3.90 × 10-6 C charge, and (c) both of the charges in (a) and (b).
Answer #1

according to gauss law

electric flux = q(enclosed) / eo

a) electric flux = 2.5*10^-6 / 8.85*10^-12

= 2.824*10^5 Nm^2/C

b) electric flux = -3.9*10^-6 / 8.85*10^-12 = -4.4*10^5 Nm^2/C

c) electric flux = (2.5 - 3.9)*10^-6 / 8.85*10^-12 = -1.58*10^5 Nm^2/C